What is the seventh term of a sequence whose first term is 1 and whose common ratio is 3?

Found 2 solutions by CubeyThePenguin, ikleyn:
Answer by CubeyThePenguin(3113)

geometric series format: a, ar, ar^2, ...

a = first term
r = common ratio

seventh term = ar^6 = (1)(3^6) = 729

Answer by ikleyn(52949)

.

On geometric progressions,  see introductory lessons
    - Geometric progressions
    - The proofs of the formulas for geometric progressions
    - Problems on geometric progressions
    - Word problems on geometric progressions
in this site.

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic
"Geometric progressions".

Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.